# Imitation Game: Reproducing Deep Learning Bugs Leveraging an Intelligent Agent

### v1

@@ Abstract.
Deep learning bugs, debugging, bug reproduction, Agentic AI
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: IEEE/ACM 48th International Conference on Software Engineering (ICSE); Apr 12–18, 2018; Rio De Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2018/06ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
To resolve Deep Learning (DL) bugs, software developers must first systematically reproduce them on their local machines. Without a reproduction, they cannot confirm the presence of a bug or diagnose its root cause. However, reproduction of DL bugs can be effort-intensive, time-consuming, and frustrating due to various technical challenges. They include intricate data pipelines, hardware dependencies, and variations in software frameworks and library versions. Even when a bug is reproducible, developers frequently need to engage in trial-and-error, carefully tune environmental settings, and reason about the contextual factors that may influence the behaviour of DL programs, all of which can be tedious and error-prone (Naziri et al., 2025). Developers also face the challenge of missing or incomplete information when attempting to reproduce bugs from issue reports (Wang et al., 2024a). Reports may lack crucial details of a bug and omit relevant data or code snippets. In such cases, even experienced developers must spend substantial time reconstructing the missing detail of a bug, the target environment, and iteratively testing hypotheses to reproduce the erroneous behaviour (Huang et al., 2023; Zhang et al., 2020). These difficulties make the manual process of reproducing DL bugs highly inefficient and error-prone. Thus, there is a strong need for techniques that can accelerate and systematize the reproduction of DL bugs, reducing developer effort while increasing the reliability of DL systems.
@@ 3.1.2. Code Snippet Retrieval
(b) ANN search: Approximate Nearest Neighbour (ANN) search leverages semantic proximity between two items within the semantic space. We use an ANN search to capture the conceptual or semantic relationship between code chunks and bug reports that may not be evident from their contents. Using a pre-trained code-language model [^2] http://bit.ly/3IwN0kk, we encode each bug report and code chunk into dense vectors and normalize them. Then, to measure the similarity between the bug report (q) and code (d), we compute their angular similarity (ang, [n. d.]).
@@ 3.1.3. Context Construction
where $D_{i}$ refers to an individual code snippet, and ResolveModule is the function that assigns it to a specific module, $M_{k}$. In this step, we cluster the code snippets according to their respective modules and determine each module’s priority based on its members’ maximum score from the re-ranking step. In other words, a module with a single highly relevant snippet will be prioritized over one with numerous less relevant snippets. After determining the module priority, we select the top k modules (e.g., 5) for subsequent analysis.
@@ 3.2. Bug Report Restructuring
First, we analyze the report title and opening paragraphs to extract the core problem statement about a bug. Second, we capture the technical details, including stack traces, error messages, performance metrics, and resource utilization patterns, to form the Observed Behaviour section. Third, we also capture the intended system behaviour from report text, documentation references, API specifications, and baseline metrics, and construct a detailed Expected Behaviour section. Finally, RepGen synthesizes all gathered information into a detailed sequence of steps that specifies environment configurations, hardware requirements, dataset specifications, and training parameters, and constructs the steps for reproducing the bug. If any steps are explicitly mentioned in a bug report, they are leveraged by the LLM. Otherwise, the LLM reasons through the gathered information and generates a set of candidate steps. Then all three items are used by the LLM to restructure and enhance the bug report.
@@ 4.1. Dataset Construction
To construct the dataset for our experiment, we followed a systematic approach inspired by prior works (Morovati et al., 2023). We select GitHub, the largest resource for open-source repositories, and use its GitHub Search API (git, [n. d.]) for repository selection. We first limit our search to Python repositories due to its frequent use in DL systems (pyt, [n. d.]). Next, we identify the repositories built on three frameworks: TensorFlow, Keras, and PyTorch, the most widely used DL frameworks (fra, [n. d.]). We collected 223,485 repositories from GitHub after this step. Then, we make sure that each repository has at least one push to the main branch after January 2023, indicating their recent maintenance (Morovati et al., 2023). After this filtering, we retained 83,649 repositories for the subsequent analysis. Finally, we select such repositories that have a minimum of 10 watchers and 500 stars, a criterion used in existing literature (Nikeghbal et al., 2023) to exclude obsolete or low-quality repositories. After applying these criteria, we collected 653 repositories from GitHub. After the preliminary filtering above, we removed repositories with fewer than 10 closed bugs, ensuring that the selected repositories have a sufficient history of reported and resolved bugs. After this filtering step, we retained 392 repositories for manual analysis. Since our primary focus was on bugs in DL systems, we manually analyzed the 392 repositories and excluded non-DL, toy or educational repositories, leading to a total of 54 repositories for subsequent analysis. We then collect all the bugs that were resolved from each of the 54 repositories in the last two years (April 1 2023 - April 1 2025). We also carefully examined each bug report, fix commits, and necessary instructions and attempted to reproduce the bug in an isolated environment. If we could reproduce the bug within 1 hour, we would include this bug in the dataset. Bugs which could not be reproduced within an hour were marked as non-reproducible and excluded from further analysis. Through this process, we identified 106 reproducible bugs from 16 distinct repositories. We spent $\approx$190 person-hours on dataset construction.
@@ 4.3. Baseline Selection
To rigorously validate our approach, RepGen, we systematically chose eight open-source and closed-source LLMs as our baseline methods. We chose Llama-3 (8B, 70B), which are open-source, decoder-only transformer models with strong instruction-following capabilities and widely adopted by the research community. We also included DeepSeek-R1 (7B, 685B) and Qwen3 (8B) for their advanced reasoning capabilities. Furthermore, we included Qwen2.5-7B, as it was the base model for RepGen. Finally, we chose GPT-4.1, a highly capable proprietary model with its long context window and advanced code generation abilities. We employed three distinct prompting techniques for every LLM (zero-shot, few-shot, and chain-of-thought), and used DSPy to optimize our prompts. For smaller models (up to 8B), we conducted experiments on a local machine using Ollama, while for larger models (70B and above), we leveraged their respective APIs (ope, [n. d.]; dee, [n. d.]; lla, [n. d.]). We generated the code using the recommended generation parameters (e.g., temperature, top_p, top_k) for the local models. To mitigate the effects of LLM non-determinism, we generated the code five times and evaluated each generation following our verification protocol (Section 4.2).
@@ 4.6.1. Answering RQ1 – Effectiveness
Statistical Significance Tests: To demonstrate RepGen’s effectiveness, we conducted paired comparisons between RepGen and each of the 24 baseline configurations using McNemar’s test on the $2\times 2$ contingency tables for the 106 shared bugs. McNemar’s test (McNemar, 1947) is appropriate for paired binary outcomes and for assessing whether the proportion of bugs reproduced by RepGen differs significantly from each baseline counterpart. For each comparison, we report the $p$-values and the Haldane–Anscombe adjusted odds ratio (hereafter, “odds ratio”) (Haldane, 1956), defined as $\mathrm{OR}=(b+0.5)/(c+0.5)$, where $b$ is the number of bugs reproduced by RepGen but not by the baseline, and $c$ is the number of bugs reproduced by the baseline but missed by RepGen. To account for multiple comparisons, we applied the Benjamini–Hochberg correction (Benjamini and Hochberg, 1995) to these tests, yielding BH-adjusted $p$-values (q-values). All 24 comparisons remained statistically significant after BH correction ($q\leq 3.80\times 10^{-3}$), with odds ratios ranging from $2.45$ to $19.86$. For example, when comparing RepGen to DeepSeek-R1-685B (few-shot), there were $b=35$ cases where RepGen succeeded and the baseline failed, versus $c=14$ cases in the opposite direction, giving an odds ratio of $2.45$ and a BH-adjusted q-value of $3.80\times 10^{-3}$. We also calculated 95% confidence intervals for the difference in mean success rates for bug reproduction over five runs between RepGen and the baseline using the Newcombe hybrid score method (Newcombe, 1998), finding that the difference is $19.8\%$ with a 95% CI of $7.0\%$$29.7\%$, indicating that RepGen reproduces between 7% and 30% more of the shared bugs than DeepSeek-R1-685B. Note that this statistic differs from the odds ratio, which summarizes the relative odds of reproduction. Overall, these results demonstrate that RepGen outperforms all baselines in bug reproduction with statistical significance.
@@ 4.6.2. Answering RQ2 – Ablation Study
C1=ANN, C2=BM25, C3=Dependency Extraction, C4=Module Partitioning, C5=Reranker, C6=Training Loop Extraction, C7=Training Loop Ranking, =Component Present
We also analyze the impact of the seven components used in our context construction step (Step 1a, Fig. 3), with results summarized in Table 3. Removal of the training loop ranking component caused the largest performance drop, reducing the success rate to 20.75%. This underscores its critical role in identifying the training loop most relevant to the reported bug, ensuring that the constructed context focuses on the correct portion of the codebase. Similarly, omitting module partitioning reduced the success rate to 23.58%, highlighting that prioritizing modules (e.g., components or classes) based on the relevance of the code snippets rather than the sheer number of snippets is essential for building an informative context. Our hybrid retrieval mechanism also proved vital for context construction. Disabling BM25 search, which captures exact lexical matches, reduced RepGen’s success rate to 24.53%, while removing ANN search, responsible for retrieving semantically related code, resulted in a 26.42% success rate. These drops demonstrate that both lexical and semantic retrieval are complementary and necessary for constructing a comprehensive context. Removal of the training loop extraction component lowered RepGen’s performance to 28.30%, indicating that incomplete loops without core training logic can weaken the context. Removing reranking, which refines the relevance of retrieved snippets, decreased success to 30.19%, showing that careful prioritization of candidate code is important for context quality. Finally, omitting dependency extraction, responsible for capturing imports, helper functions, and referenced variables, reduced RepGen’s success rate to 34.91%, highlighting the need for a complete and executable context. These findings collectively show that each element of our context construction pipeline is vital for the high performance of RepGen.
Our findings above align with earlier work (Shah et al., 2025), suggesting that successful bug reproduction requires a multi-faceted approach. The strong performance of RepGen (80.19%) compared to its ablated versions also suggests that each component addresses a specific challenge and contributes significantly when reproducing DL bugs.
@@
Data-Dependent Bug An error that is triggered by issues involving the input data, such as missing files, incorrect paths, or improper formatting. A model training script crashes during the first epoch with a NotFoundError because it cannot locate a file (17e00d... .jpg) that is listed in the dataset but is not present at the file path.
@@ 4.6.3. Answering RQ3 – Failure Analysis
To understand difficulty of reproducing certain deep learning bugs, we manually analyzed the 21 bugs that RepGen failed to reproduce. Two researchers, the lead author and an independent collaborator, each with over five years of experience in deep learning and software engineering, labelled these bugs using multiple established DL bug taxonomies (Humbatova et al., 2020; Islam et al., 2019; Morovati et al., 2024) and identified their bug types, symptoms, and root causes. We also examined each non-reproduced bug’s report and associated developer discussion to extract its symptoms and root causes, and mapped them to the closest taxonomy-defined type. The inter-rater agreement, measured using Cohen’s kappa was 0.849, which indicates an almost perfect agreement (McHugh, 2012). All disagreements were subsequently resolved through discussion between the two researchers. Based on the labelled types, we then grouped the bugs into three higher-level categories: (a) API/Dependency bugs (10/21), (b) environment-dependent bugs (5/21) and (c) data-dependent bugs (6/21), as shown in Table. 4. Reproducing these categories of bugs is challenging primarily due to their reliance on specific external conditions. For example, RepGen struggles with newer APIs and updates since they were released after the knowledge cutoff of the base model, Qwen2.5-Coder-7B. Environment-dependent bugs are also difficult to reproduce because of complex distributed systems, resource allocation across multiple GPUs, and precise hardware requirements. Similarly, data-dependent bugs are challenging due to reliance on specific datasets, file system structures, and particular data formats and paths. Moreover, we also found that each failure type aligns with distinctive pipeline signals in RepGen. API or dependency mismatches result in low BM25 or ANN relevance scores and unresolved dependency graphs. Environment-dependent bugs exhibit incomplete context coverage or unresolvable hardware-specific requirements, as indicated by runtime feedback showing that the bug cannot be reproduced in the environment created by the generated code. Data-dependent bugs produce early failures in static analysis or at runtime due to missing datasets or paths. These consistent signals explain why some bugs cannot be reproduced and provide future directions for improving RepGen’s workflow.
@@
Given the evidence above, model size is a factor, but not the sole determinant of success in reproducing DL bugs. This finding reinforces that the primary challenge of the automated DL bug reproduction is the lack of awareness of the bug’s external context. While a larger model can resolve more issues based on its broader knowledge, it still fails when the reproduction requires precise environmental, dependency, or data-related context that it cannot access. Thus, future approaches should focus on improving the quality of the context provided to LLM agents, rather than relying on increased model sizes.

### v2

@@ Abstract.
Deep learning bugs, deep learning bug reproduction, automated debugging, LLM-powered agents, code generation, machine learning systems, software testing and debugging
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3744916.3787795isbn: 979-8-4007-2025-3/2026/04ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
To resolve DL bugs, software developers must first systematically reproduce them on their local machines. Without a reproduction, they cannot confirm the presence of a bug or diagnose its root cause. However, reproduction of DL bugs can be effort-intensive, time-consuming, and frustrating due to various technical challenges. They include intricate data pipelines, hardware dependencies, and variations in software frameworks and library versions. Even when a bug is reproducible, developers frequently need to engage in trial-and-error, carefully tune environmental settings, and reason about the contextual factors that may influence the behaviour of DL programs, all of which can be tedious and error-prone (Naziri et al., 2025). Developers also face the challenge of missing or incomplete information when attempting to reproduce bugs from issue reports (Wang et al., 2024a). Reports may lack crucial details of a bug and omit relevant data or code snippets. In such cases, even experienced developers must spend substantial time reconstructing the missing detail of a bug, the target environment, and iteratively testing hypotheses to reproduce the erroneous behaviour (Huang et al., 2023; Zhang et al., 2020). These difficulties make the manual process of reproducing DL bugs highly inefficient and error-prone. Thus, there is a strong need for techniques that can accelerate and systematize the reproduction of DL bugs, reducing developer effort while increasing the reliability of DL systems.
@@ 3.1.2. Code Snippet Retrieval
(b) ANN search: Approximate Nearest Neighbour (ANN) search leverages semantic proximity between two items within the semantic space. We use an ANN search to capture the conceptual or semantic relationship between code chunks and bug reports that may not be evident from their contents. Using a code-language model [^2] http://bit.ly/3IwN0kk, we encode each bug report and code chunk into dense vectors and normalize them. Then, to measure the similarity between the bug report (q) and code (d), we compute their angular similarity (ang, 2017).
@@ 3.1.3. Context Construction
where $D_{i}$ refers to an individual code snippet, and ResolveModule is the function that assigns it to a specific module, $M_{k}$. In this step, we cluster the code snippets according to their respective modules and determine each module’s priority based on its members’ maximum score from the re-ranking step. In other words, a module with a single highly relevant snippet will be prioritized over one with numerous less relevant snippets. After determining the module priority, we select the top five modules for subsequent analysis.
@@ 3.2. Bug Report Restructuring
First, we analyze the report title and opening paragraphs to extract the core problem statement about a bug. Second, we capture the technical details, including stack traces, error messages, performance metrics, and resource utilization patterns, to form the Observed Behaviour section. Third, we also capture the intended system behaviour from report text, API specifications, and baseline metrics, and construct a detailed Expected Behaviour section. Finally, RepGen synthesizes all gathered information into a detailed sequence of steps that specifies environment configurations, hardware requirements, dataset specifications, and training parameters, and constructs the steps for reproducing the bug. If any steps are explicitly mentioned in a bug report, they are leveraged by the LLM. Otherwise, the LLM reasons through the gathered information and generates a set of candidate steps. Then all three items are used by the LLM to restructure and enhance the bug report.
@@ 4.1. Dataset Construction
To construct the dataset for our experiment, we followed a systematic approach inspired by prior works (Morovati et al., 2023). We select GitHub, the largest resource for open-source repositories, and use its GitHub Search API (git, 2022) for repository selection. We first limit our search to Python repositories due to its frequent use in DL systems (pyt, 2024). Next, we identify the repositories built on three frameworks: TensorFlow, Keras, and PyTorch, the most widely used DL frameworks (fra, 2024). We collected 223,485 repositories from GitHub after this step. Then, we make sure that each repository has at least one push to the main branch after January 2023, indicating their recent maintenance (Morovati et al., 2023). After this filtering, we retained 83,649 repositories for the subsequent analysis. Finally, we select such repositories that have a minimum of 10 watchers and 500 stars, a criterion used in existing literature (Nikeghbal et al., 2023) to exclude obsolete or low-quality repositories. After applying these criteria, we collected 653 repositories from GitHub. After the preliminary filtering above, we removed repositories with fewer than 10 closed bugs, ensuring that the selected repositories have a sufficient history of reported and resolved bugs. After this filtering step, we retained 392 repositories for manual analysis. Since our primary focus was on bugs in DL systems, we manually analyzed the 392 repositories and excluded non-DL, toy or educational repositories, leading to a total of 54 repositories for subsequent analysis. We then collect all the bugs that were resolved from each of the 54 repositories in the last two years (April 1 2023 - April 1 2025). We also carefully examined each bug report, fix commits, and necessary instructions and attempted to reproduce the bug in an isolated environment. If we could reproduce the bug within 1 hour, we would include this bug in the dataset. Bugs which could not be reproduced within an hour were marked as non-reproducible and excluded from further analysis. Through this process, we identified 106 reproducible bugs from 16 distinct repositories. We spent $\approx$190 person-hours on dataset construction.
@@ 4.3. Baseline Selection
To rigorously validate our approach, RepGen, we systematically chose eight open-source and closed-source LLMs as our baseline methods. We chose Llama-3 (8B, 70B), which are open-source, decoder-only transformer models with strong instruction-following capabilities and widely adopted by the research community. We also included DeepSeek-R1 (7B, 685B) and Qwen3 (8B) for their advanced reasoning capabilities. Furthermore, we included Qwen2.5-7B, as it was the base model for RepGen. Finally, we chose GPT-4.1, a highly capable proprietary model with its long context window and advanced code generation abilities. We employed three distinct prompting techniques for every LLM (zero-shot, few-shot, and chain-of-thought), and used DSPy to optimize our prompts. For smaller models (up to 8B), we conducted experiments on a local machine using Ollama, while for larger models (70B and above), we leveraged their respective APIs (ope, 2025; dee, 2025; lla, 2025). We generated the code using the recommended generation parameters (e.g., temperature, top_p, top_k) for the local models. To mitigate the effects of LLM non-determinism, we generated the code five times and evaluated each generation following our verification protocol (Section 4.2).
@@ 4.6.1. Answering RQ1 – Effectiveness
Statistical Significance Tests: To demonstrate RepGen’s effectiveness, we conducted paired comparisons between RepGen and each of the 24 baseline configurations using McNemar’s test on the $2\times 2$ contingency tables for the 106 shared bugs. McNemar’s test (McNemar, 1947) is appropriate for paired binary outcomes and for assessing whether the proportion of bugs reproduced by RepGen differs significantly from each baseline counterpart. For each comparison, we report the $p$-values and the Haldane–Anscombe adjusted odds ratio (hereafter, “odds ratio”) (Haldane, 1956). To account for multiple comparisons, we applied the Benjamini–Hochberg correction (Benjamini and Hochberg, 1995) to these tests, yielding BH-adjusted $p$-values (q-values). All 24 comparisons remained statistically significant after BH correction ($q\leq 3.80\times 10^{-3}$), with odds ratios ranging from $2.45$ to $19.86$. For example, when comparing RepGen to DeepSeek-R1-685B (few-shot), there were $b=35$ cases where RepGen succeeded and the baseline failed, versus $c=14$ cases in the opposite direction, giving an odds ratio of $2.45$ and a BH-adjusted q-value of $3.80\times 10^{-3}$. We also calculated 95% confidence intervals for the difference in mean success rates for bug reproduction over five runs between RepGen and the baseline using the Newcombe hybrid score method (Newcombe, 1998), finding that the difference is $19.8\%$ with a 95% CI of $7.0\%$$29.7\%$, indicating that RepGen reproduces between 7% and 30% more of the shared bugs than DeepSeek-R1-685B. Note that this statistic differs from the odds ratio, which summarizes the relative odds of reproduction. Overall, these results demonstrate that RepGen outperforms all baselines in bug reproduction with statistical significance.
@@ 4.6.2. Answering RQ2 – Ablation Study
C1=ANN, C2=BM25, C3=Dependency Extraction, C4=Module Partitioning, C5=Reranker, C6=Training Loop Extraction, C7=Training Loop Ranking
We also analyze the impact of the seven components used in our context construction step (Step 1a, Fig. 3), with results summarized in Table 3. Removal of the training loop ranking component caused the largest performance drop, reducing the success rate to 20.75%. This underscores its critical role in identifying the training loop most relevant to the reported bug, ensuring that the constructed context focuses on the correct portion of the codebase. Similarly, omitting module partitioning reduced the success rate to 23.58%, highlighting that prioritizing modules (e.g., components or classes) based on the relevance of the code snippets rather than the sheer number of snippets is essential for building an informative context. Our hybrid retrieval mechanism also proved vital for context construction. Disabling BM25 search, which captures exact lexical matches, reduced RepGen’s success rate to 24.53%, while removing ANN search, responsible for retrieving semantically related code, resulted in a 26.42% success rate. These drops demonstrate that both lexical and semantic retrieval are complementary and necessary for constructing a comprehensive context. Removal of the training loop extraction component lowered RepGen’s performance to 28.30%, indicating that incomplete loops without core training logic can weaken the context. Removing reranking, which refines the relevance of retrieved snippets, decreased success to 30.19%, showing that careful prioritization of candidate code is important for context quality. Finally, omitting dependency extraction, responsible for capturing imports, helper functions, and referenced variables, reduced RepGen’s success rate to 34.91%, highlighting the need for a complete and executable context. These findings collectively show that each element of our context construction pipeline is vital for the high performance of RepGen. Our findings above align with earlier work (Shah et al., 2025), suggesting that successful bug reproduction requires a multi-faceted approach. The strong performance of RepGen (80.19%) compared to its ablated versions also suggests that each component addresses a specific challenge and contributes significantly when reproducing DL bugs.
@@
Data-Dependent Bug An error caused by input data issues, such as missing files, incorrect paths, or improper formatting. A model training script crashes during the first epoch with a NotFoundError because it cannot locate a file (17e00d.jpg) that is listed in the dataset but is not present at the path.
@@ 4.6.3. Answering RQ3 – Failure Analysis
To understand difficulty of reproducing certain deep learning bugs, we manually analyzed the 21 bugs that RepGen failed to reproduce. Two researchers, the lead author and an independent collaborator, each with over five years of experience in deep learning and software engineering, labelled these bugs using multiple established DL bug taxonomies (Humbatova et al., 2020; Islam et al., 2019; Morovati et al., 2024) and identified their bug types, symptoms, and root causes. We also examined each non-reproduced bug’s report and associated developer discussion to extract its symptoms and root causes, and mapped them to the closest taxonomy-defined type. The inter-rater agreement, measured using Cohen’s kappa was 0.849, which indicates an almost perfect agreement (McHugh, 2012). All disagreements were subsequently resolved through discussion between the two researchers. Based on the labelled types, we then grouped the bugs into three higher-level categories: (a) API/Dependency bugs (10/21), (b) environment-dependent bugs (5/21) and (c) data-dependent bugs (6/21), as shown in Table  4. Reproducing these categories of bugs is challenging primarily due to their reliance on specific external conditions. For example, RepGen struggles with newer APIs and updates since they were released after the knowledge cutoff of the base model, Qwen2.5-Coder-7B. Environment-dependent bugs are also difficult to reproduce because of complex distributed systems, resource allocation across multiple GPUs, and precise hardware requirements. Similarly, data-dependent bugs are challenging due to reliance on specific datasets, file system structures, and particular data formats and paths. Moreover, we also found that each failure type aligns with distinctive pipeline signals in RepGen. API or dependency mismatches result in low BM25 or ANN relevance scores and unresolved dependency graphs. Environment-dependent bugs exhibit incomplete context coverage or unresolvable hardware-specific requirements, as indicated by runtime feedback showing that the bug cannot be reproduced in the environment created by the generated code. Data-dependent bugs produce early failures in static analysis or at runtime due to missing datasets or paths. These consistent signals explain why some bugs cannot be reproduced and provide future directions for improving RepGen’s workflow.
@@
Given the evidence above, model size is a factor, but not the sole determinant of success in reproducing DL bugs. This finding reinforces that the primary challenge of the automated DL bug reproduction is the lack of awareness of the bug’s external context. While a larger model can resolve more issues based on its broader knowledge, it still fails when the reproduction requires precise environment-related context that it cannot access. Thus, future approaches should focus on improving the quality of the context provided to LLM agents, rather than relying on increased model sizes.